# CAPOS: The bulge Cluster APOgee Survey VII. First detailed chemical analysis of NGC 6316

### v2

@@ Conclusions.
Globular Clusters – Spectroscopy – Abundances1 Introduction
@@
Although there is already an APOGEE pipeline that derives chemical abundances and other parameters like radial velocity for BGCs, namely the APOGEE Stellar Parameters and Chemical Abundance Pipeline (ASPCAP, García Pérez et al. (2016)), it is known that the pipeline does not perform optimally for stars with extreme types of “non-standard” elemental abundance patterns, namely so-called Second Population (2P) GC stars - those that have been enriched by some as yet-to-be-determined process (Holtzman et al., 2015; Jönsson et al., 2018). Fernández-Trincado et al. (2020) found significant differences in certain abundance ratios compared to ASPCAP, in the range of dex higher than ASPCAP values for [N, O, Al, Si, Ce/Fe] to dex lower for [O, Mg/Fe] in 2P stars. CAPOS I showed that even the ASPCAP Fe abundances of 2P stars deviates from that of their first population (1P) cousins, with the former offset to somewhat higher metallicity values within a given GC. This was strongly corroborated by Geisler et al. (2025) recent analysis of APOGEE results for all CAPOS GCs. Hence, different abundance determination techniques are necessary for 2P stars at least, and preferably one should also perform the same analysis on 1P stars in order to ensure homogeneity.
@@
According to the Baumgardt database111https://people.smp.uq.edu.au/HolgerBaumgardt/globular/ (and references therein), NGC 6316 lies at J2000 coordinates RA: 259.15542 deg and DEC: 28.14011 deg, it is located at a heliocentric distance of kpc and Galactocentric distance of 3.16 0.36 kpc, has mean absolute proper motions (PMs) of pmRA = 4.974 0.028 masyr-1 and pmDEC = 4.605 0.027 masyr-1, a mean heliocentric radial velocity (RV) of kmsec-1, a mass of M⊙, and a tidal radius (TR) pc. D23 provides a foreground reddening of E(B-V) .
Massari, Koppelman, & Helmi (2019) classify it as a Main Bulge GC, Pérez-Villegas et al. (2020) as Thick Disk, Callingham et al. (2022) as most likely originating from Kraken, and both Belokurov & Kravtsov (2024) and Chen & Gnedin (2024) as in situ. Geisler et al. (2025) give NGC 6316 an 81% probability of being an in situ GC and a 51% probability of being a BGC as opposed to Disk GC, and finally classify it as a bona fide BGC.
@@ 2 Data and target selection
The final cluster members were selected from the final version of the APOGEE-2 catalogue published in December 2021 as part of the Data Release 17 of the Sloan Digital Sky Survey (SDSS) (Abdurro’uf et al., 2022), which is available in the SDSS IV Science Archive Server222https://www.sdss4.org/dr17/irspec/spectro_data/.
@@
- Finally, based on the minimum SNR suggested by APOGEE (SNR 70)333https://www.sdss4.org/dr17/irspec/abundances/, 10 out of the 19 potential targets were included in the final membership selection.
@@
Table 1 lists the member parameters. These selections led to the previously mentioned total of 10 targets, of which 8 were observed as part of the CAPOS survey (Geisler et al., 2021), and the remaining 2 were observed under the main APOGEE survey (Majewski et al., 2017). The spectra of these 10 targets were downloaded from the SDSS IV Science Archive Server444https://data.sdss.org/sas/dr17/apogee/spectro/aspcap/dr17/synspec_rev1/. All the obtained spectra had already undergone reduction using the APOGEE-2 data reduction pipeline as described in Nidever et al. (2015).
@@ 4.1 Results and comparison with other in-situ clusters
We report abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce. These are shown in Table 2 together with overall [/Fe] abundances for each star determined with the elements Si and Ca, weighted equally. Each abundance has its associated weighted error, determined by the square root of the quadratic sum of each individual error and dividing the result by the number of stars. We adopted the solar reference abundances from Asplund, Grevesse, & Sauval (2005) for all elements, as used in most CAPOS papers, due to their widespread adoption and consistency in the literature. Table 2 also contains the micro-turbulence velocities obtained by BACCHUS. This is the first spectroscopic report of abundances manually determined for this cluster and, since we used high resolution spectroscopy, we believe the most robust study to the date. We obtain a mean metallicity for the cluster of [Fe/H] = 0.87 0.02, with no evidence of any intrinsic Fe abundance variation. Our metallicity is in very good agreement with that obtained by D23, [Fe/H] = , and Conroy et al. (2018), [Fe/H] = . However, it is some 0.5 dex lower than the canonical value of 0.45 from the H10 catalogue and similarly offset from the Vásquez et al. (2018) values of , and . As mentioned, these values were obtained using the calcium triplet method. The metallicity reported in H10 is a weighted average of the values derived by AZ88 (scaled to C09), VFO07 and C09 (triple-weighted). The AZ88 value was obtained by averaging their calcium triplet [Fe/H] estimate with that from ZW84. C09, in turn, recalibrated the ZW84 scale to derive their own metallicities. ZW84 itself is based on a transformation of the Q39 index, derived primarily from the CaII K lines, among other indicators. Thus, the high metallicities reported for this cluster by these studies likely reflect issues related to the calcium-based methods. Our metallicity determination should therefore be preferred over previous estimates since it is also based on a large sample of stars observed at high resolution and SNR. This is further supported by the fact that our study - along with that of D23 - is one of the only investigations conducted exclusively on this cluster, providing a level of detail and accuracy that is not achievable when working on large samples of GCs.
With our new metallicity, NGC 6316 is now no longer a member of the classic metal-rich BGCs. Indeed, based on both the Bica et al. (2024) and Geisler et al. (2025) metallicity distribution functions of BGCs, which (assuming that it is indeed a BGC) are bimodal with peaks near -0.5 and -1.1, NGC 6316 falls in the lower metallicity domain, which is now known to be the dominant mode.
@@
Figure 3 shows abundance relations for light elements involved in the MP phenomenon. Alongside NGC 6316 (white stars) there are four other clusters: NGC 6522 in emerald inverted triangles ([Fe/H] = , Fernández-Trincado et al. 2019), NGC 6558 in blue squares ([Fe/H] = , González-Díaz et al. 2023), Ton 2 in red triangles ([Fe/H] = , Fernández-Trincado et al. 2021) and NGC 6121 (M4) in Fuchsia circles ([Fe/H] = , Mészáros et al. 2020. These clusters were selected because all of their abundances were determined using BACCHUS, and they present similar metallicities as our analysed stars. These criteria were selected in an attempt to find similar (or different) behaviours at similar/different location for in-situ GCs. According to Geisler et al. (2025), all 5 of these GCs are in-situ, NGC 6316, NGC 6522 and NGC 6558 are indeed from the Bulge, Ton 2 is from the Disk, and it is uncertain whether NGC 6121 is from the Bulge or Disk.
@@ 4.2 The in-situ nature of NGC 6316
As mentioned, Massari, Koppelman, & Helmi (2019) and Geisler et al. (2025) classify NGC 6316 as an in-situ GC. Belokurov & Kravtsov (2024, hereafter BK24) proposed an abundance-based classification scheme for in-situ and ex-situ GCs, utilizing [Mg/Fe] vs [Al/Fe] abundances derived from ASPCAP data and other sources scaled to ASPCAP (see their Fig. 11, right panel). We replicated this scheme using all GCs from Mészáros et al. (2020) with more than five stars, systematically shifting the data following the BK24 methodology (see their Appendix A) to align the [Mg/Fe] and [Al/Fe] abundances with the BK24 separation line distinguishing in-situ and ex-situ GCs (Figure 5). In there, semi-transparent circles represent individual stars from Mészáros et al. (2020), while diamonds indicate the mean abundance for each cluster. The colour-dashed lines outline the distribution contours of individual stars. The upper and right subplots display histograms of the overall dataset’s [Mg/Fe] and [Al/Fe] distributions. The single-star distributions exhibit two distinct dense regions, corresponding to the concentrations of in-situ and ex-situ stars. However, this distinction is less clear in the mean values (diamond symbols), where the two groups overlap within a broader area.
@@
However, given the concentration of in- and ex- situ individual stars, it seems that a possible division can be made using a subgroup of stars, most probably 1P stars, that do not suffer from abundance variations. This subject, will be further explored in greater detail in a forthcoming paper (Geisler et al., 2025) focusing on 1P stars from all CAPOS clusters.
@@ Acknowledgements.
The authors acknowledge the referee’s comments that substantially improved the content of this paper. H.F. gratefully acknowledges the support provided by ALMA-ANID funds, No. 31230013. D.G.D. gratefully acknowledges the support provided by ALMA-ANID funds, No. 31230017. S.V. gratefully acknowledges the support provided by Fondecyt regular n. 1220264 and by the ANID BASAL projects FB210003. D.G. gratefully acknowledges the support provided by Fondecyt regular n. 1220264. D.G. also acknowledges financial support from the Dirección de Investigación y Desarrollo de la Universidad de La Serena through the Programa de Incentivo a la Investigación de Académicos (PIA-DIDULS). C.M. thanks the support provided by ANID GEMINI Postdoctorado No. 32230017. The team acknowledges J.G. Fernandez-Trincado for his help and commentaries. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Barrera et al. (2025) Barrera N., Villanova S., Geisler D., 2025, A&A, accepted.
@@
- Piotto et al. (2015) Piotto, G., Milone, A. P., Bedin, L. R., et al. 2015, AJ, 149, 91
@@
- Saito et al. (2024) Saito R. K., Hempel M., Alonso-García J., Lucas P. W., Minniti D., Alonso S., Baravalle L., et al., 2024, arXiv, arXiv:2406.16646. doi:10.48550/arXiv.2406.16646
@@ Appendix A Comparisons with ASPCAP
Given the substantial number of new abundances obtained, it is a good opportunity to compare them with the ASPCAP values provided in the APOGEE value-added catalogue from Schiavon et al. (2024, hereafter S24). A similar analysis, expanded for all CAPOS cluster, is being carried out in Geisler et al. (2025). S24 presented abundances for a group of stars in NGC 6316, which includes all 10 of our target stars, enabling a reliable comparison. For brevity, we will refer to our results as BACCHUS and those from S24 as ASPCAP.

### v3

@@ Abstract
Galaxy: bulge – globular clusters: individual: NGC 6316 – Techniques: spectroscopic – Stars: abundances1 Introduction
@@
Although there is already an APOGEE pipeline that derives chemical abundances and other parameters like radial velocity for BGCs, namely the APOGEE Stellar Parameters and Chemical Abundance Pipeline (ASPCAP, García Pérez et al. (2016)), it is known that the pipeline does not perform optimally for stars with extreme types of “non-standard” elemental abundance patterns, namely so-called Second Population (2P) GC stars - those that have been enriched by some as yet-to-be-determined process (Holtzman et al., 2015; Jönsson et al., 2018). Fernández-Trincado et al. (2020) found significant differences in certain abundance ratios compared to ASPCAP, in the range of dex higher than ASPCAP values for [N, O, Al, Si, Ce/Fe] to dex lower for [O, Mg/Fe] in 2P stars. CAPOS I showed that even the ASPCAP Fe abundances of 2P stars deviates from that of their first population (1P) cousins, with the former offset to somewhat higher metallicity values within a given GC. This was strongly corroborated by Geisler et al. (in prep) recent analysis of APOGEE results for all CAPOS GCs. Hence, different abundance determination techniques are necessary for 2P stars at least, and preferably one should also perform the same analysis on 1P stars in order to ensure homogeneity.
@@
According to the Baumgardt database222https://people.smp.uq.edu.au/HolgerBaumgardt/globular/ (and references therein), NGC 6316 lies at J2000 coordinates RA: 259.15542 deg and DEC: 28.14011 deg, it is located at a heliocentric distance of kpc and Galactocentric distance of 3.16 0.36 kpc, has mean absolute proper motions (PMs) of pmRA = 4.974 0.028 masyr-1 and pmDEC = 4.605 0.027 masyr-1, a mean heliocentric radial velocity (RV) of kmsec-1, a mass of M⊙, and a tidal radius (TR) pc. D23 provides a foreground reddening of E(B-V) .
Massari, Koppelman, & Helmi (2019) classify it as a Main Bulge GC, Pérez-Villegas et al. (2020) as Thick Disk, Callingham et al. (2022) as most likely originating from Kraken, and both Belokurov & Kravtsov (2024) and Chen & Gnedin (2024) as in situ. Geisler et al. (in prep) give NGC 6316 an 81% probability of being an in situ GC and a 51% probability of being a BGC as opposed to Disk GC, and finally classify it as a bona fide BGC.
@@ 2 Data and target selection
The final cluster members were selected from the final version of the APOGEE-2 catalogue published in December 2021 as part of the Data Release 17 of the Sloan Digital Sky Survey (SDSS) (Abdurro’uf et al., 2022), which is available in the SDSS IV Science Archive Server333https://www.sdss4.org/dr17/irspec/spectro_data/.
@@
- Finally, based on the minimum SNR suggested by APOGEE (SNR 70)444https://www.sdss4.org/dr17/irspec/abundances/, 10 out of the 19 potential targets were included in the final membership selection.
@@
Table 1 lists the member parameters. These selections led to the previously mentioned total of 10 targets, of which 8 were observed as part of the CAPOS survey (Geisler et al., 2021), and the remaining 2 were observed under the main APOGEE survey (Majewski et al., 2017). The spectra of these 10 targets were downloaded from the SDSS IV Science Archive Server555https://data.sdss.org/sas/dr17/apogee/spectro/aspcap/dr17/synspec_rev1/. All the obtained spectra had already undergone reduction using the APOGEE-2 data reduction pipeline as described in Nidever et al. (2015).
@@ 4.1 Results and comparison with other in-situ clusters
We report abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni, and Ce. These are shown in Table 2 together with overall [/Fe] abundances for each star determined with the elements Si and Ca, weighted equally. Each abundance has its associated weighted error, determined by the square root of the quadratic sum of each individual error and dividing the result by the number of stars. We adopted the solar reference abundances from Asplund, Grevesse, & Sauval (2005) for all elements, as used in most CAPOS papers, due to their widespread adoption and consistency in the literature. Table 2 also contains the micro-turbulence velocities obtained by BACCHUS. This is the first spectroscopic report of abundances manually determined for this cluster and, since we used high-resolution spectroscopy, we believe the most robust study to date. We obtain a mean metallicity for the cluster of [Fe/H] = 0.87 0.02, with no evidence of any intrinsic Fe abundance variation. Our metallicity is in very good agreement with that obtained by D23, [Fe/H] = , and Conroy et al. (2018), [Fe/H] = . However, it is some 0.5 dex lower than the canonical value of 0.45 from the H10 catalogue (that is a weighted average of the values derived by AZ88 scaled to C09, VFO07, and C09. The latter triple-weighted) and similarly offset from the Vásquez et al. (2018) values of , and . As mentioned, these values were obtained by relatively uncertain photometric or low resolution spectroscopic means, and/or using a smaller number of samples. We also cannot discard uncertainties in the membership of some stars in their samples. Thus, our metallicity determination should therefore be preferred over previous estimates, since it is based on a large sample of stars with a high probability of membership, observed at high resolution and SNR.
With our new metallicity, NGC 6316 is now no longer a member of the classic metal-rich BGCs. Indeed, based on both the Bica et al. (2024) and Geisler et al. (in prep) metallicity distribution functions of BGCs, which (assuming that it is indeed a BGC) are bimodal with peaks near -0.5 and -1.1, NGC 6316 falls in the lower metallicity domain, which is now known to be the dominant mode.
@@
Figure 3 shows abundance relations for light elements involved in the MP phenomenon. Alongside NGC 6316 (white stars) there are four other clusters: NGC 6522 in emerald inverted triangles ([Fe/H] = , Fernández-Trincado et al. 2019), NGC 6558 in blue squares ([Fe/H] = , González-Díaz et al. 2023), Ton 2 in red triangles ([Fe/H] = , Fernández-Trincado et al. 2021) and NGC 6121 (M4) in Fuchsia circles ([Fe/H] = , Mészáros et al. 2020. These clusters were selected because all of their abundances were determined using BACCHUS, and they present similar metallicities as our analysed stars. These criteria were selected in an attempt to find similar (or different) behaviours at similar/different location for in-situ GCs. According to Geisler et al. (in prep), all 5 of these GCs are in-situ, NGC 6316, NGC 6522 and NGC 6558 are indeed from the Bulge, Ton 2 is from the Disk, and it is uncertain whether NGC 6121 is from the Bulge or Disk.
@@ 4.2 The in-situ nature of NGC 6316
As mentioned, Massari, Koppelman, & Helmi (2019) and Geisler et al. (in prep) classify NGC 6316 as an in-situ GC. Belokurov & Kravtsov (2024, hereafter BK24) proposed an abundance-based classification scheme for in-situ and ex-situ GCs, utilizing [Mg/Fe] vs [Al/Fe] abundances derived from ASPCAP data and other sources scaled to ASPCAP (see their Fig. 11, right panel). We replicated this scheme using all GCs from Mészáros et al. (2020) with more than five stars, systematically shifting the data following the BK24 methodology (see their Appendix A) to align the [Mg/Fe] and [Al/Fe] abundances with the BK24 separation line distinguishing in-situ and ex-situ GCs (Figure 5). In there, semi-transparent circles represent individual stars from Mészáros et al. (2020), while diamonds indicate the mean abundance for each cluster. The colour-dashed lines outline the distribution contours of individual stars. The upper and right subplots display histograms of the overall dataset’s [Mg/Fe] and [Al/Fe] distributions. The single-star distributions exhibit two distinct dense regions, corresponding to the concentrations of in-situ and ex-situ stars. However, this distinction is less clear in the mean values (diamond symbols), where the two groups overlap within a broader area.
@@
However, given the concentration of in- and ex- situ individual stars, it seems that a possible division can be made using a subgroup of stars, most probably 1P stars, that do not suffer from abundance variations. This subject, will be further explored in greater detail in a forthcoming paper (Geisler et al. in prep) focusing on 1P stars from all CAPOS clusters.
@@ Acknowledgements.
The authors acknowledge the referee’s comments that substantially improved the content of this paper. H.F. gratefully acknowledges the support provided by ALMA-ANID funds, No. 31230013. D.G.D. gratefully acknowledges the support provided by ALMA-ANID funds, No. 31230017. S.V. gratefully acknowledges the support provided by Fondecyt regular n. 1220264 and by the ANID BASAL projects FB210003. D.G. gratefully acknowledges the support provided by Fondecyt regular n. 1220264. D.G. also acknowledges financial support from the Dirección de Investigación y Desarrollo de la Universidad de La Serena through the Programa de Incentivo a la Investigación de Académicos (PIA-DIDULS). C.M. thanks the support provided by ANID GEMINI Postdoctorado No. 32230017. The team acknowledges J.G. Fernandez-Trincado for his help and commentaries. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Barrera et al. (2025) Barrera N., Villanova S., Geisler D., Fernández-Trincado J. G., Muñoz C., 2025, A&A, 699, A128. doi:10.1051/0004-6361/202451968
@@
- Piotto et al. (2015) Piotto G., Milone A. P., Bedin L. R., Anderson J., King I. R., Marino A. F., Nardiello D., et al., 2015, AJ, 149, 91. doi:10.1088/0004-6256/149/3/91
@@
- Saito et al. (2024) Saito R. K., Hempel M., Alonso-García J., Lucas P. W., Minniti D., Alonso S., Baravalle L., et al., 2024, A&A, 689, A148. doi:10.1051/0004-6361/202450584
@@ Appendix A Comparisons with ASPCAP
Given the substantial number of new abundances obtained, it is a good opportunity to compare them with the ASPCAP values provided in the APOGEE value-added catalogue from Schiavon et al. (2024, hereafter S24). A similar analysis, expanded for all CAPOS cluster, is being carried out in Geisler et al. (in prep). S24 presented abundances for a group of stars in NGC 6316, which includes all 10 of our target stars, enabling a reliable comparison. For brevity, we will refer to our results as BACCHUS and those from S24 as ASPCAP.